# Bridging Voting and Deliberation with Algorithms: Field Insights from vTaiwan and Kultur Komitee

### v1

@@ Abstract.
Democratic processes increasingly aim to integrate large-scale voting with face-to-face deliberation, addressing the challenge of reconciling individual preferences with collective decision-making. This work introduces new methods that use algorithms and computational tools to bridge online voting with face-to-face deliberation, tested in two real-world scenarios: Kultur Komitee 2024 (KK24) and vTaiwan. These case studies highlight the practical applications and impacts of the proposed methods. We present three key contributions: (1) Radial Clustering for Preference Based Subgroups, which enables both in-depth and broad discussions in deliberative settings by computing homogeneous and heterogeneous group compositions with balanced and adjustable group sizes; (2) Human-in-the-loop MES, a practical method that enhances the Method of Equal Shares (MES) algorithm with real-time digital feedback. This builds algorithmic trust by giving participants full control over how much decision-making is delegated to the voting aggregation algorithm as compared to deliberation; and (3) the ReadTheRoom deliberation method, which uses opinion space mapping to identify agreement and divergence, along with spectrum-based preference visualisation to track opinion shifts during deliberation. This approach enhances transparency by clarifying collective sentiment and fosters collaboration by encouraging participants to engage constructively with differing perspectives. By introducing these actionable frameworks, this research extends in-person deliberation with scalable digital methods that address the complexities of modern decision-making in participatory processes.
@@ 1. Introduction
Fair mechanisms for aggregating diverse preferences are essential for proportional representation in collective decision-making, ensuring outcomes reflect the priorities of all groups rather than only favouring majorities (Aziz and Shah, 2020). Algorithms like the Method of Equal Shares (MES) (Peters et al., 2021) enable fair resource allocation while protecting minority voices. However, real-world implementation faces challenges, including bridging technological literacy gaps, managing real-time deliberations, and building trust in algorithmic outcomes. To address these issues, accessible and user-friendly tools are vital for translating theoretical fairness into practical, inclusive applications across diverse contexts (Yang et al., 2024).
This research offers practical insights into designing democratic innovations that integrate voting and deliberation. In particular, we introduce three computational methods applied in two real-world settings: Kultur Komitee Winterthur 2024 (KK24) and vTaiwan. The contributions of this paper are:
We introduce Radial Clustering for Preference Based Subgroups, which computes homogeneous or heterogeneous subgroups for deliberation based on prior voting results. This clustering method ensures that subgroups have balanced sizes by slicing the (two-dimensional) preference space into even parts.
We extend the Method of Equal Shares (MES) with a Human-in-the-Loop approach, allowing participants to adjust the overall budget and explore funding scenarios, balancing algorithmic decision-making with deliberation.
We present the ReadTheRoom Deliberation method, which uses opinion mapping to identify divisive statements and spectrum-based voting to visualise shifts in preferences. This feedback loop fosters reflection, collaboration, and open-minded discussions.
We document two participatory case studies: KK24 and vTaiwan. KK24 demonstrates the impact of a Budget Assembly for inclusive funding decisions, while vTaiwan shows how combining online and offline deliberation enhances participation and mutual learning.
@@ 1.1. Case Studies
1.1.1. Kultur Komitee: Democratising Art and Culture through Participatory Budgeting in Winterthur, Switzerland
@@
Out of the 37 participants involved in the process, 35 attended the final deliberative workshop (), with a mean age of 41.2 years, ranging from 20 to 82. The gender distribution consisted of 13 males and 22 females. Most participants were born in Switzerland, while 8 were from outside Switzerland. By the end of the process, 35 projects were funded, receiving grants ranging from CHF 2,500 to CHF 30,000, totalling CHF 381,500. During the process of Human-in-the-loop MES, it was decided that 50% of the budget would be used for MES calculation, and the other half would be allocated with deliberation. Eventually, 18 projects were selected using MES, and 17 projects through deliberation. KK24 serves as a model for integrating participatory and deliberative practices in resource allocation. For more details, we refer to the KK24 factsheet 111https://kulturkomitee.win/media/digital_kk23_24_factsheet.pdf.
@@
Launched in 2014, vTaiwan is a decentralised open consultation process combining online and offline interactions to connect citizens and government for the deliberation on national issues. It serves as a model to involve government ministries, elected representatives, scholars, experts, business leaders, civil society organisations, and citizens in crafting digital legislation. One of vTaiwan’s key tools is Polis (Small et al., 2021), a digital platform for opinion collection and consensus formation. The tool facilitates large-scale conversations and has been pivotal in achieving “rough consensus” on various policy issues at the national level, overcoming scalability challenges in deliberative democracy. Since its inception, vTaiwan has engaged over 200,000 participants and contributed to 26 pieces of legislation on the digital economy and other key issues.
In December 2024, vTaiwan and the Taiwan Network Information Center (TWNIC) co-hosted a Roundtable Discussion on “What aspects should Taiwan consider when regulating artificial intelligence (AI)?” This event addressed legal and ethical challenges of generative AI and gathered input for Taiwan’s proposed Basic Law on Artificial Intelligence. Using Polis and Mentimeter, public sentiment was visualised, areas of consensus and disagreement were highlighted, and opinion shifts were tracked during workshops. This deliberative process provided actionable insights for AI governance while showcasing the effectiveness of participatory approaches to address complex policy challenges.
@@ 2.1. Hybrid Democratic Innovation
Hybrid Democratic Innovations (HDIs) combine concentrated deliberation with large-scale voting (Hendriks and Michels, 2024) to enhance inclusivity, reflection, and policy impact. Early examples of HDIs were given by Fishkin (1991) with deliberative polling. This method has been described by Mansbridge (2010) to capture informed public opinion through random sampling, structured deliberation, and feedback. Building on this, Felicetti (2021) proposed “democratic assemblage,” emphasising adaptive systemic integration of deliberative and participatory processes. This has been illustrated by Gastil and Broghammer (2021), who showed how digital platforms and game mechanics can enhance participatory budgeting through motivation and engagement. Practical applications include the three-step model of Itten and Mouter (2022) for climate policymaking, which combines mini-publics and maxi-publics to produce widely accepted recommendations, and Hendriks and Michels (2024)’s study of Antwerp’s Citizens’ Budget, showcasing how “participatory budgeting new-style” can align deliberation with large-scale voting to uphold democratic values. Referendum processes have similarly benefited, as Witting et al. (2025) identify how deliberation mitigates deficits of referenda. Hendriks and Wagenaar (2023) propose deliberative referenda to connect citizens’ assemblies with direct voting, thereby enhancing public debate and systemic legitimacy. Finally, Chambers and Warren (2023) theorise the complementary role of voting in deliberation to provide closure, equality, and accountability, highlighting the importance of hybrid approaches in advancing democratic innovation.
@@ 2.2. Online Tools for Deliberative Democracy
Digital tools have emerged as critical innovations in deliberative democracy, enabling scalable citizen participation in decision-making (Klein, 2017; Davies and Procter, 2020; Mikhaylovskaya, 2024). These tools integrate deliberation with digital methods for visualising, moderating, and summarising discussions. A prominent example is Polis (Small et al., 2021), an open-source platform where participants collect and rate statements on given topics. The tool visualises the high-dimensional opinion space and computes consensus statements to foster common ground in the deliberation. Being applied to many larger participatory processes such as vTaiwan and the German “Aufstehen” [“Stand up”] movement, Polis demonstrates how digital tools can scale deliberation beyond traditional in-person settings. Another successful example of such an online tool is the Stanford Deliberation Platform (Fishkin et al., 2019), where automatic moderation ensures fair deliberation of small groups in video calls. This platform has also been used in the “America in One Room” deliberation. To summarise discussions, tools like Wikum (Zhang et al., 2017) and Kialo (Chaudoin et al., 2017) have been developed. These tools offer thread-like discussions with features such as nested argumentation and can be used to challenge reasoning in deliberation. Decision-making platforms such as Decidim (Aragón et al., 2017) facilitate participatory processes for cities, while initiatives like Decide Madrid and vTaiwan exemplify algorithmic empowerment, decentralising power and fostering pluralistic policymaking (Tseng, 2022). This emphasises the possible role of algorithms in enhancing democratic transparency and inclusivity. Umbelino et al. (2021) introduce ProtoTeams, a gamified approach to forming diverse groups for real-time collaborative discussions, illustrating how team composition and dynamics influence decision-making outcomes. However, scaling deliberation introduces challenges, as many platforms prioritise technical solutions, often overlooking factors like cultural and linguistic diversity, social inequalities, and inclusion across genders and abilities (Shortall et al., 2022). Addressing these aspects is also essential to ensure digital tools support a broader and more diverse range of participants.
2.3. Computational Social Choice, Voting and Participatory Budgeting
@@
Computational social choice provides theoretical and practical tools for designing voting systems that address fairness, proportionality, and collective decision-making in contexts such as participatory budgeting (PB) (Brandt et al., 2016). Fairness in this context refers to the equal representation of diverse voter preferences and the proportional allocation of resources or decision-making power, ensuring that no group is disproportionately advantaged or excluded (Fain et al., 2016). Recent advancements have emphasised the integration of computational methods with participatory processes. Aziz and Shah (2020) provide a comprehensive survey of PB models, highlighting the importance of preference elicitation, voter incentives, and welfare objectives. Furthermore, the subfield of multi-winner voting contributes insights into proportionality and robustness. Brill and Peters (2023) proposes robust axioms for proportional representation, ensuring fairness for cohesive voter groups. Faliszewski et al. (2023) further demonstrate the utility of proportional rules in PB elections. Peters et al. (2021) formalises axioms for proportional representation in PB with additive utilities, proposing rules like the Method of Equal Shares (MES) to achieve fairness and efficiency. The basic idea of the Method of Equal Shares is that each voter is assigned an equal part of the budget. This part of the budget can only be used to fund projects that the voter has voted for. The method goes through all project proposals, beginning with the projects with the highest number of votes. It selects a project if it can be funded using the budget shares of those who voted for the project. The method then divides the cost of a project equally among its supporters. Faliszewski et al. (2023) demonstrates how MES achieves fairer outcomes and avoids divisive districts while improving overall utility. Additionally, Yang et al. (2024) explore digital PB platforms, identifying design principles that reduce cognitive load while enhancing fairness and legitimacy, where participants favour expressive input formats and fair budget allocations. Building on these contributions, our paper explores how combining deliberation and computational tools for democratic decision-making should be designed to achieve proportional fairness, transparency, and practicality.
@@ 3. Methodology
This study employs a mixed methods approach, combining computational techniques, digital tools, and participatory frameworks to explore the integration of voting and deliberation. Our methodology is structured around three algorithmic methods: Radial Clustering for Preference Based Subgroups, Human-in-the-loop MES, and ReadTheRoom Deliberation. As these methods have been co-designed and co-developed specifically to address the real-world issues and needs of KK24 and vTaiwan, they are also applied and evaluated in the context of these two case studies.
3.1. Radial Clustering for Preference Based Subgroups
@@ 3.1.1. Motivation
Group deliberation often leads to middle-ground solutions that fail to address niche interests, such as local or specific community needs. This tendency for deliberation to gravitate towards consensus solutions has been noted in the literature, with scholars arguing that structured approaches are needed to ensure the representation of marginalised or niche perspectives (Abdullah et al., 2016; Karpowitz and Raphael, 2014; Sunstein, 2017). Fraser (1990) highlights the importance of creating spaces for marginalised groups to deliberate autonomously, while Mansbridge (1994) emphasises the role of “enclaves of protected discourse” to empower these groups in broader discussions. Sunstein (2017) introduces the term “Enclave Deliberation,” noting both its potential to address inequalities and the risks of group polarisation, which can be mitigated through careful integration into wider deliberative processes.
Radial Clustering for Preference Based Subgroups, hereby referred to as just “Radial Clustering”, builds on the concept of Enclave Deliberation by structuring initial discussions in homogeneous groups. In these groups, participants with similar voting patterns deliberate to ensure niche interests are well-represented and given a platform. This stage creates an environment where marginalised perspectives can emerge without being overshadowed by dominant views. In the second round, participants transition to heterogeneous groups to engage with diverse perspectives, building broader understanding and identifying projects that benefit larger segments of citizens. By combining enclave-inspired homogeneous groups with broader heterogeneous deliberation, Radial Clustering aim to ensure both equality and inclusivity in the decision-making process.
Standard clustering methods, such as k-Means (MacQueen et al., 1967), often create uneven groups or exclude outliers. While extensions exist that adjust the loss function to balance cluster sizes (Malinen and Fränti, 2014), these complex algorithms might not be very comprehensive for participants in the deliberation. Our suggestion, Radial clustering, ensures equal group sizes with a straightforward and visually understandable algorithm: It projects the participants to a two-dimensional opinion space and divides it radially into “pizza slices”. Despite not being the most optimal clustering mathematically, this algorithm achieves balanced assignments and ensures visually transparent explainability, making it practical for real-world settings.
3.1.2. Process
@@
In KK24, the organising committee decided that six groups with 6-7 participants each were needed for group deliberations. The steps for group assignment and deliberation were as follows:
@@
Dimensionality Reduction and Mapping: The group composition was based on participants’ voting similarity in a pre-survey (i.e., approval votes for 50+ projects). These high-dimensional data were reduced using Principal Component Analysis (PCA), projecting participants’ preferences into a two-dimensional space. Each participant’s angular position in this space was calculated relative to the mean position:
where and represent the coordinates of participant . The angles were normalised to 0-360 degrees.
@@
Radial Partitioning and Group Assignment: The angular space was divided into six sectors, each having 6-7 participants. To achieve this, each sector was initialised spanning degrees. Participants were assigned to sectors based on their angular positions. The angles of each sector were then iteratively adjusted to arrive at balanced group sizes.
Two-Round Deliberation with Homogeneous and Heterogeneous Groups: The Radial clustering method was used to assign participants to six homogeneous groups (A/B/C/D/E/F) in the first round. In the second round, participants were reassigned to six heterogeneous groups (1/2/3/4/5/6), ensuring diversity by mixing members from different homogeneous groups.
@@
In both rounds, group members discussed, selected, and ranked five projects. Rankings were converted into Borda scores (e.g., 1st: 5 points, 2nd: 4 points, etc.), and the points from all groups were aggregated across both rounds. Projects were then selected based on total points until the budget was fully allocated.
3.1.3. Evaluation Metric
To evaluate the alignment between individual votes and group decisions, we calculate a weighted, normalised alignment score () for each voter in both homogeneous and heterogeneous rounds. For a voter , the score is defined as , where is the set of projects the voter supported and were also selected by the group, is the weight assigned to project , and is the total number of projects the voter supported. This calculation is performed separately for the homogeneous and heterogeneous rounds to obtain two distinct alignment scores for each voter.
3.2. Human-in-the-Loop Method of Equal Shares (Human-in-the-Loop MES)

### v2

@@ Abstract.
Democratic processes increasingly integrate large-scale voting with face-to-face deliberation to reconcile individual preferences with collective decision-making. This work introduces algorithmic methods to bridge online voting with face-to-face deliberation, tested in two real-world scenarios: Kultur Komitee 2024 (KK24) and vTaiwan. We present three key contributions: (1) Preference-based Clustering for Deliberation (PCD), enabling both focused and broad discussions by computing balanced homogeneous and heterogeneous groups; (2) Human-in-the-loop MES, enhancing the Method of Equal Shares algorithm with real-time feedback, giving participants control over algorithmic decision-making; and (3) the ReadTheRoom method, using opinion mapping to identify agreement and divergence while tracking opinion shifts during deliberation. These actionable frameworks extend in-person deliberation with scalable digital methods that address the complexities of modern participatory decision-making.
111© The definitive Version of Record was published in:
Proceedings of the 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greece.
DOI: https://doi.org/10.1145/3715275.3732205
@@ 1. Introduction
Fair mechanisms for aggregating diverse preferences are essential for proportional representation in collective decision-making, ensuring outcomes reflect the priorities of all groups rather than only favouring majorities (Aziz and Shah, 2020). Algorithms like the Method of Equal Shares (MES) (Peters et al., 2021) enable fair resource allocation while protecting minority voices. However, real-world implementation faces challenges, including bridging technological literacy gaps, managing real-time deliberations, and building trust in algorithmic outcomes. To address these issues, accessible and user-friendly tools are vital for translating theoretical fairness into practical, inclusive applications (Yang et al., 2024b).
This research offers practical insights into designing democratic innovations that integrate voting and deliberation. In particular, we introduce three computational methods applied in two real-world settings: Kultur Komitee Winterthur 2024 (KK24) and vTaiwan. The contributions of this paper are: (1) We introduce Preference-based Clustering for Deliberation (PCD), a proposed framework that utilises voting data to create more effective deliberation groups. The process focuses on using opinion clustering to group participants of similar and dissimilar preferences together to create different deliberation dynamics. (2) We extend the Method of Equal Shares (MES) with a Human-in-the-Loop approach, allowing participants to adjust the overall budget and explore funding scenarios, balancing algorithmic decision-making with deliberation. (3) We present the ReadTheRoom deliberation method, which uses opinion mapping to identify divisive statements and spectrum-based voting to visualise shifts in preferences. This feedback loop fosters reflection, collaboration, and open-minded discussions. (4) We document two participatory case studies. KK24 demonstrates the impact of a Budget Assembly for inclusive funding decisions, while vTaiwan shows how combining online and offline deliberation enhances participation and mutual learning.
@@ 1.1. Case Studies
1.1.1. Kultur Komitee: Democratising Art and Culture through Budgeting Assembly
@@
Out of the 37 participants involved in the process, 35 attended the final deliberative workshop (), with a mean age of 41.2 years, ranging from 20 to 82. The gender distribution consisted of 13 males and 22 females. Most participants were born in Switzerland, while 8 were from outside Switzerland. By the end of the process, 35 projects were funded, receiving grants ranging from CHF 2,500 to CHF 30,000, totalling CHF 381,500. During the process of Human-in-the-loop MES, it was decided that 50% of the budget would be used for MES calculation, and the other half would be allocated with deliberation. Eventually, 18 projects were selected using MES, and 17 projects through deliberation. KK24 serves as a model for integrating participatory and deliberative practices in resource allocation. The authors were invited to implement the MES calculation and co-design the process with the organising committee. For more details, we refer to the KK24 factsheet 222https://kulturkomitee.win/media/digital_kk23_24_factsheet.pdf.
@@
Launched in 2014, vTaiwan is a decentralised open consultation process combining online and offline interactions to connect citizens and government for deliberation on national issues. It serves as a model for involving diverse stakeholders in crafting digital legislation, using tools like Polis (Small et al., 2021) to facilitate large-scale conversations and achieve consensus. Since inception, vTaiwan has engaged over 200,000 participants and contributed to 26 pieces of legislation. In December 2024, vTaiwan and the Taiwan Network Information Center co-hosted a roundtable on AI regulation, gathering input for Taiwan’s proposed Basic Law on Artificial Intelligence. Using Polis and Mentimeter, the event visualised public sentiment, highlighted areas of agreement and disagreement, and tracked opinion shifts during deliberation. The corresponding author offered the ReadTheRoom Deliberation method concept through participation in vTaiwan’s meetings but primarily remained as observers, analysing only anonymised open data published through vTaiwan’s public repositories. This dual role as both community member and researcher is acknowledged as providing valuable insider perspective while potentially introducing bias, addressed through transparent methodology and reliance on publicly available data.
@@ 2.1. Hybrid Democratic Innovation
Hybrid Democratic Innovations (HDIs) combine concentrated deliberation with large-scale voting (Hendriks and Michels, 2024) to enhance inclusivity and policy impact. Fishkin (1991) pioneered this approach with deliberative polling, which Mansbridge (2010) describes as capturing informed public opinion through random sampling and structured deliberation. Felicetti (2021) expanded this concept with “democratic assemblage,” emphasising systemic integration of deliberative and participatory processes. Gastil and Broghammer (2021) demonstrated how digital platforms enhance participatory budgeting through engagement. Practical applications include Itten and Mouter (2022)’s three-step climate policymaking model combining mini-publics and maxi-publics, and Hendriks and Michels (2024)’s analysis of Antwerp’s Citizens’ Budget showing how “participatory budgeting new-style” aligns deliberation with voting. For referendums, Witting et al. (2025) show how deliberation addresses process deficits, while Hendriks and Wagenaar (2023) propose deliberative referenda connecting citizens’ assemblies with direct voting. Finally, Chambers and Warren (2023) theorises how voting complements deliberation by providing closure, equality, and accountability, highlighting the importance of hybrid approaches.
@@ 2.2. Online Tools for Deliberative Democracy
Digital tools have emerged as critical innovations in deliberative democracy, enabling scalable citizen participation in decision-making (Klein, 2017; Davies and Procter, 2020; Mikhaylovskaya, 2024). A prominent example is Polis (Small et al., 2021), an open-source platform where participants rate statements on given topics. The tool visualises opinion spaces and computes consensus statements to foster common ground in deliberation. Used in initiatives like vTaiwan and the German “Aufstehen” movement, Polis demonstrates how digital tools can scale deliberation beyond traditional settings. Similarly, the Stanford Deliberation Platform (Fishkin et al., 2019) uses automatic moderation for fair small-group video deliberations, as seen in the “America in One Room” project. For discussion summarisation, tools like Wikum (Zhang et al., 2017) and Kialo (Chaudoin et al., 2017) offer nested argumentation features. Decision-making platforms such as Decidim (Aragón et al., 2017) facilitate city-level participation, while Decide Madrid and vTaiwan exemplify algorithmic empowerment for pluralistic policymaking (Tseng, 2022). Umbelino et al. (2021)’s ProtoTeams demonstrates how team composition influences decision-making outcomes through gamified group formation. However, scaling deliberation presents challenges, as many platforms prioritise technical solutions while overlooking cultural and linguistic diversity, social inequalities, and inclusion across different demographics (Shortall et al., 2022). Addressing these aspects is essential for supporting a broader range of participants.
2.3. Voting and Participatory Budgeting
@@
Computational social choice provides tools for designing voting systems that address fairness, proportionality, and collective decision-making in participatory budgeting (PB) (Brandt et al., 2016). Fairness here means equal representation of diverse preferences and proportional resource allocation (Fain et al., 2016). Aziz and Shah (2020) survey PB models, highlighting preference elicitation, voter incentives, and welfare objectives. In multi-winner voting, Brill and Peters (2023) proposes robust axioms for proportional representation, while Faliszewski et al. (2023) demonstrates proportional rules’ utility in PB elections. Peters et al. (2021) formalises proportional representation axioms in PB, proposing the Method of Equal Shares (MES) for fairness and efficiency. With MES, each voter receives an equal budget share usable only for projects they support. The method evaluates projects by vote count, selecting those fundable using supporters’ budget shares, then dividing costs equally among those supporters. Faliszewski et al. (2023) shows MES achieves fairer outcomes and improves utility. In Yang et al. (2024b)’s experiment, participants perceive MES as significantly fairer than conventional Greedy voting. Our paper explores how combining deliberation with digital tools for democratic decision-making can achieve proportional fairness, transparency, and practicality.
@@ 3. Methodology
This study employs a mixed methods approach, combining computational techniques, digital tools, and participatory frameworks to explore the integration of voting and deliberation. Our methodology is structured around three algorithmic methods: Preference-based Clustering for Deliberation, Human-in-the-loop MES, and ReadTheRoom Deliberation. As these methods have been co-designed and co-developed specifically to address the real-world issues and needs of KK24 and vTaiwan, they are also applied and evaluated in the context of these two case studies. Also, the real-world implementation of these democratic innovations required pragmatic evaluation approaches that could capture both objective outcomes and participant experiences. For each method, we developed context-specific metrics that balanced theoretical rigour with practical measurement constraints in field settings.
3.1. Preference-based Clustering for Deliberation (PCD)
@@ 3.1.1. Motivation
Traditional deliberative processes typically treat voting and deliberation as separate mechanisms rather than complementary parts of a unified decision-making system. This separation means valuable information from prior voting is often underutilised when forming discussion groups. The proposed Preference-based Clustering for Deliberation (PCD) framework bridges this gap by using voting data to create more effective deliberation groups.
@@
Group deliberation often leads to middle-ground solutions that fail to address niche interests, such as local or specific community needs. This tendency for deliberation to gravitate towards consensus solutions has been noted in the literature, with scholars arguing that structured approaches are needed to ensure the representation of marginalised or niche perspectives (Abdullah et al., 2016; Karpowitz and Raphael, 2014; Sunstein, 2017).
The theoretical foundations for this approach draw from several key concepts in deliberative democracy. Fraser (1990) highlights the importance of creating spaces for marginalised groups to deliberate autonomously, while Mansbridge (1994) emphasises the role of “enclaves of protected discourse” to empower these groups in broader discussions. Sunstein (2017) introduces the term “Enclave Deliberation,” noting both its potential to address inequalities and the risks of group polarisation, which can be mitigated through careful integration into wider deliberative processes.
3.1.2. Process and Implementation
The PCD framework implements these ideas through a three-phase process: Initially, all participants cast their votes on the projects, proposals or opinions under consideration. These voting patterns are then used to inform the subsequent deliberation phases. The deliberation unfolds in two rounds:
(1) First, participants with similar voting patterns are grouped together, creating homogeneous deliberation spaces where niche interests can be fully explored and articulated without being overshadowed by majority preferences.
(2) Then, participants are reconfigured into diverse groups with different voting patterns, allowing for heterogeneous deliberation that builds broader consensus while ensuring minority viewpoints have already been developed. This balanced approach ensures both the depth of preference exploration and the breadth of inclusive decision-making across the deliberative process.
While the PCD framework can be implemented using various clustering algorithms, for the KK24 case study we employed a Radial Clustering method. This represents just one of many possible approaches to operationalise the framework’s core principle. However, most standard clustering algorithms like k-Means (MacQueen et al., 1967) create uneven groups or exclude outliers. While balanced variants exist (Malinen and Fränti, 2014), they can be difficult for non-technical facilitators to work with. The Radial Clustering method addresses these challenges by projecting participants into a two-dimensional opinion space and dividing it into “pizza slice” sectors (see Figure 1). Each sector contains an equal number of participants, ensuring balanced group sizes.
To demonstrate this, we evaluated six clustering approaches using the KK24 pre-deliberation voting data, as detailed in Table 3 in the appendix. The results show trade-offs between mathematical optimality and practical usability. Radial Clustering achieves a reasonable silhouette score (0.238) while maintaining perfectly balanced groups and clear visual separation between clusters. This visual simplicity offers practical advantages in real-world settings—when participants cannot attend the deliberation event, facilitators can quickly adjust the sector boundaries on printed diagrams. The broader PCD framework remains implementation agnostic, organisations can select the specific clustering algorithm that best suits their technical capabilities and deliberative context. The essential innovation lies in the PCD framework, systematically incorporating voting information to enhance deliberative processes and creating a more integrated approach to collective decision-making.
3.1.3. Process with Radial Clustering
@@
In KK24, the organising committee determined that six groups with 6-7 participants each were needed for effective deliberation. We implemented Radial Clustering through the following steps:
@@
(1) Dimensionality Reduction and Mapping: Group composition was based on participants’ voting patterns from the pre-deliberation survey (approval votes for 50+ projects). Using Principal Component Analysis (PCA), we reduced this high-dimensional data to a two-dimensional space and calculated each participant’s angular position relative to the mean using , where and represent the coordinates of participant . The angles were normalised to 0-360 degrees.
@@
(2) Radial Partitioning and Group Assignment: We divided the angular space into six sectors. Each sector was initially spanning degrees. Participants were assigned to sectors based on their angular positions, then sector boundaries were iteratively adjusted to ensure balanced group sizes of 6-7 people per group.
@@
(3) Two-Round Deliberation: We conducted two deliberation rounds with different group compositions. For the first round, participants were placed in six homogeneous groups (A/B/C/D/E/F) with others sharing similar voting patterns. In the second round, they were reconfigured into six heterogeneous groups (1/2/3/4/5/6), ensuring diversity by mixing members from different homogeneous groups. In both rounds, groups discussed, selected and ranked five projects, with rankings converted to Borda scores (e.g., 1st: 5 points, 2nd: 4 points). Projects were selected based on aggregate points from all groups across both rounds until the budget was fully allocated.
3.1.4. Evaluation Metrics
To evaluate the PCD framework, we measured: preference alignment (), quantifying how well group decisions reflected individual preferences. This alignment was calculated as the proportion of projects that both appeared in a participant’s individual votes and received points from their deliberation group; voting-deliberation correlation, analysing statistical relationships between pre-deliberation votes and group decisions; resource allocation patterns, comparing project cost distributions between rounds using Mann-Whitney U Tests; demographic effects, examining alignment differences across gender and age groups with Wilcoxon signed-rank tests; and participant experience, using Likert-scale ratings and qualitative feedback on decision-making ease and preference representation. This approach captured both objective outcomes and subjective experiences across different group compositions.
3.2. Human-in-the-Loop Method of Equal Shares